OpenStudy (anonymous):

2x^4+2x^3-5x^2-4x+1 divided by x+14 By the Remainder Theorem, if P(x) is divided by x + 14 = x −(?) The remainder is P (?) The question marks are where the answers should go...im not quite sure what its asking because i did the division and the R is 80613 but it says it is incorrect.

OpenStudy (anonymous):

http://www.webmath.com/polydiv.html

OpenStudy (anonymous):

We are using synthetic division.

OpenStudy (anonymous):

the remainder is P(-14) P(x) = Q(x) * D(x) + R(x) if D(x) = x+14 then if x = -14, D(x) =0 and P(-14) = R(-14) which is the remainder (since D(x) is a first degree and the remainder will have a degree less than that of D(x), meaning the degree of R(x) will be 0, meaning R(x) will be a constant).

OpenStudy (anonymous):

also, x+14 = x - (?)... what value of x makes the left side equal 0? that is what the ? equals.

OpenStudy (anonymous):

the ? in the parentheses.